OpenStudy (anonymous):

I WILL GIVE MEDAL just help me please what are the solutions to this equation|x-10|-4=2x

Vocaloid (vocaloid):

the absolute value sign means we need to solve for two cases, one where (x-10) is positive and one where (x-10) is negative so we get: 1. x - 10 - 4 = 2x 2. -(x-10) - 4 = 2x solve each equation for x

OpenStudy (anonymous):

1. x - 10 - 4 = 2x do i add 4 with -4 and 2x?

Vocaloid (vocaloid):

add 4 to both sides of the equal sign and tell me what you get

OpenStudy (anonymous):

1. x - 10 - 4 = 2x +4 +4 x-10=2x+4 is it like that?

Vocaloid (vocaloid):

yes, now keep going

OpenStudy (anonymous):

Why not x - 10 - 4 = 2x -> x - 14 = 2x ?

Vocaloid (vocaloid):

^ that works, too

OpenStudy (anonymous):

Then subtract x from both sides

OpenStudy (campbell_st):

don't worry about that equation as it doesn't have a valid solution solve the other equation

OpenStudy (anonymous):

im still confuse with the first one..

Vocaloid (vocaloid):

ok, we can pick up where you left off x-10=2x+4 now add 10 to both sides

OpenStudy (anonymous):

x=2x+14 is it like that?

Vocaloid (vocaloid):

right, now subtract 2x from both sides

OpenStudy (anonymous):

x=2x+14 -2x -2x x=12x

Vocaloid (vocaloid):

not quite, remember that you can't combine 14 since there's no x attached to it x - 2x = ?

OpenStudy (anonymous):

oh ok so its x-2x=14-2x

Vocaloid (vocaloid):

almost x-2x = 2x+14-2x now keep going x - 2x = ?

OpenStudy (anonymous):

12

Vocaloid (vocaloid):

nope 1-2 = ?

OpenStudy (anonymous):

-1

Vocaloid (vocaloid):

right, so if 1-2 = -1, what does x - 2x = ?

OpenStudy (anonymous):

-x

Vocaloid (vocaloid):

correcto, now we have -x = 2x+14-2x what is 2x+14-2x?

OpenStudy (anonymous):

im not sure

Vocaloid (vocaloid):

hint: 2x - 2x = ?

OpenStudy (anonymous):

0

Vocaloid (vocaloid):

good, so what is 2x+14-2x?

OpenStudy (anonymous):

0

Vocaloid (vocaloid):

not quite... 2x+14-2x = 2x-2x+14 = ?

OpenStudy (anonymous):

It may be easier if it's re written as 2x - 2x + 14 = ?

OpenStudy (anonymous):

@Vocaloid never mind I didn't realize you wrote the same thing

OpenStudy (anonymous):

12

Vocaloid (vocaloid):

nope, not quite.. 2x - 2x = 0 2x - 2x + 14 = ?

OpenStudy (anonymous):

2x - 2x + 14 => 0 + 14

OpenStudy (anonymous):

14

Vocaloid (vocaloid):

right, that gives us -x = 14, which makes x = -14. (this solution doesn't quite work, like campbell mentioned, we'll see why later) now on to the second equation ~

Vocaloid (vocaloid):

-(x-10) - 4 = 2x

OpenStudy (anonymous):

-(x-10) - 4 = 2x -x+10-4=2x

Vocaloid (vocaloid):

good, keep going 10-4 = ?

OpenStudy (anonymous):

6

Vocaloid (vocaloid):

good, that gives us -x + 6 = 2x now add x to each side

OpenStudy (anonymous):

-x + 6 = 2x +x +x -x+6x+2x+x

Vocaloid (vocaloid):

not quite, remember that you can't combine numbers and variables -x + x + 6 = 2x + x

Vocaloid (vocaloid):

now keep going -x + x + 6 = ?

OpenStudy (anonymous):

6

Vocaloid (vocaloid):

excellent! now, x + 2x = ?

OpenStudy (anonymous):

2x

Vocaloid (vocaloid):

nope, here's a hint: 1 + 2 = 3 x + 2x = ?

OpenStudy (anonymous):

3x

Vocaloid (vocaloid):

good, now we have 3x = 6 now divide both sides by 3

OpenStudy (anonymous):

x=2

Vocaloid (vocaloid):

good, we're almost done after solving the two equations we have x = -14 and x = 2 our next step is to substitute back into the original equation to see which ones work and which ones don't when we plug in x = -14, we get |x-10|-4=2x |-14-10|-4 = 2(-14) |-24|-4 = -28 24 - 4 = -28 20 = -28 (this is not true, so x = -14 is not our solution)

Vocaloid (vocaloid):

next, we try x = 2 |x-10|-4=2x |2-10| -4 = 2(2) |-8| - 4 = 4 8 - 4 = 4 4 = 4 this is true, so x = 2 is our only solution

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

so it's x=-14 and x=2

OpenStudy (dm91):

i would say it is because of the work that you guys did up above but if not i will go back through it with you.